Probability: Complement of an Event

Interactive · 1 question

The Complement of an event is the case where the event does not occur. Usually we add a bar over the symbol for an event to denote its complement, so the complement of AA is Aˉ\bar{A}.

The probability of the complement of an event is P(Aˉ)=1−P(A).P(\bar{A}) = 1 - P(A). Let's say we're curious about whether it will snow tomorrow. We would say that the event of interest AA is "It snows tomorrow." Therefore, asking about P(A)P(A) means that we want to know "What is the probability that it snows?" If we ask about P(Aˉ)P(\bar{A}), then we want to know "What is the probability of it not snowing tomorrow?" If the weatherman says that there is a 36%36\% chance of snow, then by the logic of complements, there is a 64%64\% chance of not snow - which includes the possibility that the sun will shine, that it will rain or hail, and the possibility of the inevitable heat death of the universe finally occurring. Many other unlikely, but possible states of the weather tomorrow, too.

If there is a 3%3\% chance of cats and dogs raining from the sky, what is the probability of its complement (no cats and dogs falling from the stratosphere)?

Choose one answer

Interactive · 5 questions

Complements

No, these aren't compliments, but rather complements. Just in case you're disappointed, you should know that we think you're awesome!

In probability and statistics, the complement of an event is the probability that the event doesn't occur.

In other words, the complement of event A is the probability that event A does not occur.

Here's how we abbreviate the complement in statistics language:

P(A) = probability of event A
P(A') = complement event A

You might be wondering, "How the HECK am I supposed to find the probability of something not happening?"

Luckily, it's a pretty straightforward formula:

P(A') = 1 - P(A)

Let's start with a simple example.

We flip a coin and it lands on heads. What is the complement of this event?

Choose one answer

Interactive · 3 questions

Let's say I am engaged in a frivolous card game with my friends. We somehow decide to try to guess the value of a random card. At first there are two cards — a 33 and a 77. If the probability of me guessing a 33 is 33%33\%, what is the probability of me guessing a 77?

Choose one answer

Video · + 2 min read

(Video) Probability of the Complement of an Event

by HCCMathHelp

HCCMathHelp has a great video showing how to find the probability of a complement of an event.

Summary\large{\text{Summary}}

The complement of an event is the probability of an event NOT happening. If the probability of an event is P(A)P(A), then the complement of AA is Aˉ\bar{A} and therefore the probability of the complement happening is P(Aˉ)P(\bar{A}).

To calculate Aˉ\bar{A}, you only need to use this simple equation: P(Aˉ)=1−P(A)P(\bar{A})= 1- P(A)

This is because the probability of all of your events must add up to 11, or 100100%.

The first example in the video uses a question based on airline flights. "An airline reports that 9494% of its flights arrive on time, what is the probability that a flight will be late?"

For this problem we are going to call the chance that the airline is on time AA. Here, AA is 9494% or A=0.94A=0.94. Aˉ\bar{A} is probability that the plane does not arrive on time, so we can use our equation to find Aˉ\bar{A}.

P(Aˉ)=1−P(A)P(Aˉ)=1−0.94P(Aˉ)=0.06\begin{align}P(\bar{A})&= 1- P(A)\\ P(\bar{A})&= 1 - 0.94\\ P(\bar{A})&= 0.06 \end{align}

The chance that the airline is not on time is 66%.

The second example used shows a spinner with 55 sections labeled AA, BB, CC, DD, and EE. The probability of AA is P(A)=0.23P(A)=0.23. He wants to find the probability of the complement of AA, or Aˉ\bar{A}. Once again you just have to use the equation.

P(Aˉ)=1−P(A)P(Aˉ)=1−0.23P(Aˉ)=0.77\begin{align} P(\bar{A}) &= 1 - P(A)\\ P(\bar{A}) &= 1 - 0.23\\ P(\bar{A}) &= 0.77\\ \end{align}

The probability of the spinner landing on not AA is 0.770.77.

The third example is as follows: "In a bag containing 2929 marbles, 55 of the marbles are red, and 22 of the marbles are green. What is the probability of randomly selecting a marble that is not red nor green?"

First we need to find the probability of getting a red or green marble. P(R)P(R) will be the probability of getting a red marble, and P(G)P(G) will be the probability of getting a green marble.

P(R)=529P(G)=229\begin{align} P(R) &= \frac{5}{29}\\ P(G) &= \frac{2}{29}\\ \end{align}

Now we're going to add the probability of both events together to get P(R∪G)P(R\cup G)

P(R∪G)=529+229P(R∪G)=729\begin{align} P(R \cup G) &= \frac{5}{29} + \frac{2}{29}\\ P(R \cup G)&=\frac{7}{29} \end{align}

The probability of getting the complement of getting red or green, or the chance that you pull a marble that isn't red or green, is P(R∪Gˉ)P(\bar{R \cup G}). To calculate this, we can use P(R∪Gˉ)=1−P(R∪G)P(\bar{R \cup G}) = 1 - P(R \cup G)

P(R∪Gˉ)=1−P(R∪G)P(R∪Gˉ)=1−729P(R∪Gˉ)=2229\begin{align} P(\bar{R \cup G}) &= 1 - P(R \cup G)\\ P(\bar{R \cup G}) &= 1 - \frac{7}{29}\\ P(\bar{R \cup G}) &= \frac{22}{29} \end{align}

The probability of not pulling a red or green marble is 2229\frac{22}{29}.

Type a concept, like “slope” or “photosynthesis”.

↑ ↓ to choose · Enter to openSearch explanations and problems